OpenStudy (anonymous):

how much interest will you have to pay for a credit card balance of $814 that is 2 months overdue, if a 19% annual rate is charged?

OpenStudy (anonymous):

simple interest would be (2/12)(0.19)($814)

OpenStudy (ranga):

Annual interest rate = 19% = 0.19 Monthly interest rate = .19/12. We can use the compound interest formula to compute the balance, subtract the principal and find the interest. What is not clear is: credit card bills are monthly and if two payments have been skipped ("2 months overdue") are they asking for how much interest will be chnarged when the statement is received on the third month or are they asking for interest exactly one day after the second month bill was due.

OpenStudy (ranga):

Assuming two months, t = 2 Balance = Loan amount * (1 + 0.19/2)^(2) = 814 * (1 + 0.19/2)^(2) = ? Interest = Balance - Loan Amount = Balance - 840 = ?

OpenStudy (ranga):

I am getting interest at $25.98 at the end of 2 months. But if they are asking for the interest rate when the next bill comes then it will be t =3. But most probably they mean t =2 although I wish they had stated it a bit clearly.